The length of a rectangle is 6 units. The area of the rectangle is numerically equal to its perimeter. The area of the rectangle is A square units. Find A

Answer by ptaylor(2198)

Perimeter(P) of rectangle =2*Length+2*Width=2L+2W
Area(A) of rectangle=Length*Width=LW
Let W=width of rectangle
Then area, A, of rectangle=6W
And Perimeter(P), of rectangle=2*6+2W=12+2W
So, our equation to solve for the width is:
6W=12+2W subtract 2W from each side
6W-2W=12+2W-2W
4W=12
W=3 units---------------width of rectangle
A=LW=6*3=18 sq units-----------------------ans
CK
P=2*3+2*6=18
A=3*16=18
18=18
Hope this helps---ptaylor